HUBBLE SPACE TELESCOPE DAILY REPORT #1609 PERIOD COVERED: 0000Z (UTC) 04/09/96 - 0000Z (UTC) 04/10/96 Daily Status Report as of 101/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 5892 (Physical Conditions in the Galactic Halos = of the Milky Way and NGC-306) The High Resolution Spectrograph (Side Two) was used to make = observations of the radio galaxy/quasar 3C232 in order to study the metal = absorption line system produced in the line of sight by the gas in the = Milky Way and galaxy NGC-3076 halos. The observations were executed as = scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 6363 (The Environments and Host = Galaxies of BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of = BL-Lac object MS0958.9+2 in order to study its environment and its host = galaxies. The goal is to investigate systematic trends in host galaxy = type, unusual morphological features, nearby companion galaxies, etc. = The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 6250 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor) The WF/PC-2 was used to make a series on internal calibration = observations. Internal flats, biases, and K-Spot calibration images were = taken. These observations are designed to monitor the health of the = cameras. The observations were executed as scheduled, and no anomalies = were noted. 1.4 Completed Two Sets of GHRS/1 5886 (An In-Depth Study of Transition = Region Physics: Capella) The High Resolution Spectrograph (Side One) was used to make = observations of the bright star Capella (aka, HD34029) to learn more = about the transition regions that produce the ultraviolet emission lines. = A wavelength calibration was also taken. The observations were = completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Operations Notes: In preparation for the simulation being run today, HST operations = personnel moved to the SAMS backup facility at 101/0600Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: GSFC Internal Simulation #3 (EVA #3 - DIU/SSR/SADE) will be = conducted from approximately 8:00 a.m. to 8:00 p.m today. /DMH